OpenStudy (anonymous):

Which line is a graph of the equation: 2x+5y=-10 A. line a B. line b C. line c D. line d

OpenStudy (cwrw238):

first rearrange your equation to a form where the slope of the line is obvious 2x + 5y = -10 5y = -2x - 10 divide through by 5:- y = (-2/5)x - 2 this is the form y = mx + c where m = slope and c = intercept on the y-axis a negative slope is this way: \

OpenStudy (cwrw238):

now examine the 4 lines and identify the one with the slope -2/5 and intercept on y-axis = -2

OpenStudy (anonymous):

So D?

OpenStudy (cwrw238):

can you see which one it is? its a negative slope so its either b or d

OpenStudy (cwrw238):

right

OpenStudy (cwrw238):

as you can see this line passes through the point where y = -2

OpenStudy (anonymous):

Thx :D